two alcohol substances, a 15% solution and a 45% solution. How much of each solution should be used to obtain 300 milliliters of a 25% solution?
Answer by mananth(16946)

percent ---------------- quantity
Alcohol I 15 ---------------- x ml
Alcohol II 45 ------ 300 - x ml
Mixture 25.00% ---------------- 300

15x+45(300-x)=25*300

15x+13500-45x =7500
15 x - 45 x= 7500 - 13500
-30 x= -6000
/ -30
x =200 ml 15.00% Alcohol I
100 ml 45.00% Alcohol II